Operations Report – April 10th

Operations Report 10-APR-2019

SOL: n/a

Name of person filing report: Shannon Rupert

Non-nominal systems: Tunnels

Notes on non-nominal systems: About 50 % of the tunnel system has been damaged in the winds. No structural damage but most of the tarps are gone or damaged. More high winds today made repairs impossible.

Generator (hours run): Off at 8 am, on again at 7 pm

Solar— SOC 98 %. For some reason, even though it was cloudy all day, the power system worked beautifully.

Diesel Reading – 46 %

Propane Reading – 48 %

Ethanol Free Gasoline – 0 gallons.

Water (auxillary tank) – 0 gallons

Water (static tank) – 75 gallons

Auxillary to Static tank transfer– no

Gallons transferred: No

Water in GreenHab – unknown gallons.

Water (loft) – Static to Loft Pump used – no

Water Meter: not read

Toilet tank emptied: no

Deimos rover used: Still non-operational


Beginning charge:

Ending charge:

Currently charging:

Sojourner rover used: ASSIGNED TO DIRECTOR, no

Hours: 50.7

Beginning charge: 100

Ending charge: 100

Currently charging: yes

Spirit rover used: No

Hours: 86.5

Beginning charge: 100

Ending charge: 100

Currently charging: yes

Opportunity rover used: No

Hours: 56.4

Beginning charge: 100

Ending charge: 100

Currently charging: yes

Curiosity rover used: No

Hours: 88.8

Beginning charge: 100

Ending charge: 100

Currently charging: yes

Notes on rovers: Nothing to report

ATV’s Used: (Honda, 300, 350.1, 350.2, 350.3)

Reason for use: None were used.

Oil Added? No

ATV Fuel Used: 0 Gals.

# Hours the ATVs were Used today:0

Notes on ATVs: nothing to report

HabCar used and why, where? No

CrewCar used and why, where? Off-site

General notes and comments: Extreme winds again today. No MOOP was picked up.

Summary of internet: Nothing to report

Summary of suits and radios: Nothing to report

Summary of Hab operations: Front airlock still not fixed due to high winds

Summary of GreenHab operations: Still not done with the cleanup.

Summary of ScienceDome operations: Still need to know who left the biohazard containers and what is in them. (see Health and Safety)

Summary of RAM operations: Nothing to report

Summary of any observatory issues: Nothing to report

Summary of health and safety issues: Does anyone know who left the regulated medical waste at MDRS? There are two small yellow biohazard containers out by the generator that should have been removed by the crew who generated them. As this is a regulated waste, I need to know what is in it and how to dispose of it. Who left it? Please follow up with this, everyone. Thank you.

I would also like reassurances from the crew who had the experiment in the GreenHab with radishes that had people names as the labels that there was no urine or fecal matter or any other human waste (or other animal waste) in the soil. Those pots were very rotten and smelled horrible. Please follow up on this everyone and let me know. Thank you.

Questions, concerns, and requests to Mission Support:

Are the two pumps on the ground working or not working?

Why are there two batteries in the EVA room and what are they for?

See Health and Safety Summary.

Copyright © The Mars Society. All rights reserved. | Main Site